OpenStudy (anonymous):

25cm3 of a solution of O.2M potassium hydroxide reacts with 30cm3 of a solution of nitric acid. What is the concentration of the acid in g dm-3?

OpenStudy (anonymous):

Sorry, I'm late!

OpenStudy (anonymous):

what does dm stand for? Decimeter?

OpenStudy (anonymous):

Yes. Zeta, don't be apologetic.

OpenStudy (anonymous):

Ok so just use M1V1 = M2V2 25ml NaOH x 0.2M NaOH = M of HNO3 x 30ml HNO3 M of HNO3 = 0.167 M or grams / mol

OpenStudy (anonymous):

grams / Liter, not "mol" lol sorry

OpenStudy (anonymous):

so....

OpenStudy (anonymous):

1L is just 1meter ^3

OpenStudy (anonymous):

so... 0.167 mol / 1meter^3 x (10^3 decimeters / 1 meter^3) = 167 mol/decimeter

OpenStudy (anonymous):

*correciton again, the initial answer I got was 0.167 mol / L not "g/mol or g/L" LOl sorry

OpenStudy (anonymous):

so it should be 167 mol/decimeter^3

OpenStudy (anonymous):

My answer is coming to be 10.5.

OpenStudy (anonymous):

Let me rewrite this neatly

OpenStudy (anonymous):

Ok so just use M1V1 = M2V2 25ml NaOH x 0.2M NaOH = M of HNO3 x 30ml HNO3 M of HNO3 = 0.167 M or 0.167 mol/L 1L = 1 meter^3 so 0.167 mol / 1 meter^3 x ( [10 decimeter]^3 / 1 meter^3) = 167 mol / decimeter^3

OpenStudy (anonymous):

oh you want units of "grams" / decimeter ^3 ok sorry lets set that straight

OpenStudy (anonymous):

167 mol HNO3 / 1 decimeter^3 x ( 63 grams HNO3 / 1 mol HNO3) = 10521 grams / decimeter^3. I would like it more if the denominator was in meters^3 (or liters) so that the answer would be 10.521 grams / meters^3

OpenStudy (anonymous):

So ya you're right the answer is 10.5 but you can only get that if your denominator is "meters^3" NOT "DECImeter^3" as you had told me

OpenStudy (anonymous):

so with everything neatly again Ok so just use M1V1 = M2V2 25ml NaOH x 0.2M NaOH = M of HNO3 x 30ml HNO3 M of HNO3 = 0.167 M or 0.167 mol/L 1L = 1 meter^3 so 0.167 mol / 1 meter^3 x (63 grams HNO3 / 1 mol HNO3) = 10.521 grams / meter^3 of HNO3

OpenStudy (anonymous):

First of all, your reaction is the following:\[\rm \color{#C00}1KOH + \color{#C00}1HNO_3 \to \color{#C00}1KNO_3 + \color{#C00}1H_2O\]This tells you that the number of moles of both \(\rm KOH\) and \(\rm HNO_3\) are the same. The question indirectly gives you the number of moles of \(\rm KOH\). It tells you that the molarity of \(\rm KOH\) is 0.2, meaning that the number of moles of \(\rm KOH\) in \(\rm 1000 cm^3\) of solution is 0.2. Here, you have \(\rm 25 cm^3\) so the number of moles of \(\rm KOH\) would be \(\frac{25}{1000} = \frac{1}{40}\). You have the same number of moles for nitric acid (remember why?). So there are \(1/40\) moles of nitric acid which means that its mass is \(1/40 \times (1 + 14 + 48) = 1/40 \times (M_0) = M_0/40 ~\rm grams\). You are given the volume to be \(\rm 30 cm^3\). \(\rm 1 dm^3\) is \(1000 \rm cm^3\), so the volume of this solution in cubic-decimeters would be \(\frac{30}{1000} =\rm \frac{3}{100} dm^3\). Dividing the two, you will get\[\rm conc = \dfrac{M_0/40}{3/100} = \dfrac{100 /40 \times M_0}{3} = \dfrac{2.5}{3} \times 63 = 2.5 \times 21 = 52.5\]

OpenStudy (anonymous):

Oh no...

OpenStudy (anonymous):

Um, no Zeta, that's wrong

OpenStudy (anonymous):

I've already solved it

OpenStudy (anonymous):

Yes, that's wrong by a factor of five because I did not take the 0.2 into consideration. The answer would be \(52.5 \times 1/5 = 10.5\)

OpenStudy (anonymous):

I think my way is simpler haha

OpenStudy (anonymous):

Sure.

OpenStudy (anonymous):

I sense some competition here. Anyway, thank you both.

OpenStudy (anonymous):

Obviously I know more chem than Zeta haha

OpenStudy (anonymous):

We have the same approach though. :) M1V1 = M2V2 is not necessarily true, so I think we must first list down our reaction and compare like I did.

OpenStudy (anonymous):

It's definitely true b/c this is a neutralization reaction so the moles of acid must equal the moles of base. Another way or writing moles is Molarity x Volume (b/c it comes from the equation Molarity = moles / Volume). So moles acid = moles base is the same thing as Macid x Vacid = Mbase x Vbase

OpenStudy (anonymous):

But ya you can see the 1:1 mole relationship via the equation too

OpenStudy (anonymous):

\[\rm 2NaOH + H_2CO_3 \to N_2CO_3 + 2 NaOH\]

OpenStudy (anonymous):

this isn't a neutralization (which occurs between strong acid and strong base) but instead a weak acid - strong base titration

OpenStudy (anonymous):

Well, fair enough.

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

Though looking at it from a stoichiometric point of view, it may not be necessary that the coefficients are always the same. So I guess you're right... darn!!! I give up.

OpenStudy (anonymous):

lmao

OpenStudy (anonymous):

We're both right, just you did it the long way I did it the short way

OpenStudy (anonymous):

Alright I'm out, PEACE!

OpenStudy (anonymous):

That's a subjective way to think about it. I made it one long reply, plus I showed all steps of my work. So yeah.

OpenStudy (anonymous):

this is my entire final reply: 25ml NaOH x 0.2M NaOH = M of HNO3 x 30ml HNO3 M of HNO3 = 0.167 M or 0.167 mol/L 1L = 1 meter^3 so 0.167 mol / 1 meter^3 x (63 grams HNO3 / 1 mol HNO3) = 10.521 grams / meter^3 of HNO3

OpenStudy (anonymous):

that's it

OpenStudy (anonymous):

I'm pretty sure this is objectively shorter than your reply lol

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

Sorry im just bored lol

OpenStudy (anonymous):

\[\rm moles = \dfrac{25}{1000}\times \dfrac{1}{5} = \dfrac{1}{200}\]\[\rm mass = \dfrac{63}{200} g\]\[\rm volume = \dfrac{30}{1000} dm^3\]\[\rm conc = mass/volume = 10.5 gdm^{-3}\]

OpenStudy (anonymous):

You win brah lol. I'm out now. peace